# Properties of Immersions for Systems with Multiple Limit Sets with Implications to Learning Koopman Embeddings

### v1

@@ 5 Extensions of the Main Theorem
Due to the condition (T2), Theorem 3.16 cannot be directly applied to show the non-existence of one-to-one linear immersions for systems with diverging trajectories in $\mathcal{X}$. In this section, we show several extensions to Theorem 3.16 that work for systems with diverging trajectories.
5.1 Indirect Extensions of Theorem 3.16
The following two propositions, in conjunction with Theorem 3.16, show the non-existence of one-to-one linear immersions even when a diverging trajectory is present. We refer to these propositions as “indirect extensions” to Theorem 1 because they must be utilized in conjunction with it.
Proposition 5.31
@@
Proposition 5.32
Suppose that $\mathcal{X}^{\prime}\subseteq\mathcal{X}$ is forward invariant for the time-reversed system $\dot{x}=-f(x)$. If $F$ immerses the system $\dot{x}=f(x)$ over ${\mathcal{X}}$ into $\dot{z}=g(z)$ over ${\mathcal{Z}}$ by $F$, then the same $F$ also immerses the time-reversed system $\dot{x}=-f(x)$ over ${\mathcal{X}}^{\prime}$ into $\dot{z}=-g(z)$ over $F({\mathcal{X}}^{\prime})$.
@@
Thus, the solution $\psi(-t,F(\xi))$ of $\dot{z}=-g(z)$ is well defined and equal to $F(\varphi(-t,\xi))$ for all $t\geq 0$. $\Box$ By Proposition 5.32, if we can show that a one-to-one linear immersion does not exist for the time-reversed system $\dot{x}=-f(x)$ over a forward-invariant subdomain of ${\mathcal{X}}$, then there is no one-to-one linear immersion for the original system $\dot{x}=f(x)$ over ${\mathcal{X}}$. The following example demonstrates how to extend our results to systems with diverging trajectories by combining the above two propositions with Theorem 3.16.
@@
Consider a system in $\mathbb{R}^{2}$ with the phase portrait shown in Fig. 4(a). The system has three limit sets $\{-1,0,1\}$, where $-1$ and $1$ are unstable equilibria and 00 is a saddle point. We want to show that there is no one-to-one linear immersion $F$ of this system over ${\mathcal{X}}=\mathbb{R}^{2}$. To achieve this goal, we cannot directly apply Corollary 3.17 since any trajectory starting outside $[-1,1]\times{0}$ is not precompact. However, by applying Theorem 3.16 to the forward-invariant subdomain ${\mathcal{X}}^{\prime}=[-1,0]\times\{0\}$ (or $[0,1]\times\{0\}$) of ${\mathcal{X}}$, we know that there is no one-to-one linear immersion over ${\mathcal{X}}^{\prime}$. Then, by Proposition 5.31, no one-to-one linear immersion exists over ${\mathcal{X}}$.
@@
To make this example more challenging, consider ${\mathcal{X}}=\mathbb{R}^{2}\backslash\left((-1,0.5)\cup(0.5,1)\right)\times\{0\}$. Note that our previous argument does not work anymore since any forward invariant ${\mathcal{X}}^{\prime}\subseteq{\mathcal{X}}$ contains at least one trajectory that is not precompact. Now consider the time-reversed system, shown by the phase portrait in Fig. 4(b). Since ${\mathcal{X}}$ is not forward invariant for the time-reversed system, we take the forward invariant subset ${\mathcal{X}}^{\prime}=\mathbb{R}^{2}\backslash\left((-1,0)\cup(0,1)\right)\times\{0\}$ of ${\mathcal{X}}$. Since all trajectories of the time-reversed systems in ${\mathcal{X}}^{\prime}$ are precompact, Corollary 3.17 implies that there is no one-to-one linear immersion for the time-reversed system over ${\mathcal{X}}^{\prime}$. Then, by Proposition 5.32, we know there is no one-to-one linear immersion for the original system over ${\mathcal{X}}$.
5.2 Direct Extension of Theorem 3.16
@@
While Propositions 5.31 and 5.32 allow us to apply Theorem 3.16 to systems with diverging trajectories, identifying the appropriate subdomain ${\mathcal{X}}^{\prime}$ for more complex examples can be challenging. In this section, we present a direct extension to Theorem 3.16, where we replace the original condition (T2) with a weaker condition. Before delving into this extension, we introduce several key definitions.
@@
For a system $\dot{x}=f(x)$ defined over ${\mathcal{X}}$, let ${\mathcal{X}}^{\prime}$ be the maximal subset of ${\mathcal{X}}$ such that $\varphi(-t,\xi)$ is defined and contained in ${\mathcal{X}}$ for all $t\geq 0$ and $\xi\in{\mathcal{X}}^{\prime}$ (Namely ${\mathcal{X}}^{\prime}$ is the maximal forward invariant subset of ${\mathcal{X}}$ with respect to the time-reversed system). Given an initial state $\xi\in X$, the trajectory $\varphi(t,\xi)$ from $\xi$ is called backward precompact in ${\mathcal{X}}$ if $\xi\in X^{\prime}$ and the trajectory from $\xi$ is precompact in ${\mathcal{X}}$ with respect to the time-reversed system $\dot{x}=-f(x)$.
@@
For each initial state $\xi\in{\mathcal{X}}$, the $\omega$-limit set of $\xi$ in ${\mathcal{X}}$ with respect to the time-reversed system $\dot{x}=-f(x)$ is denoted by $\omega^{-}(\xi)$. If $\xi\not\in{\mathcal{X}}^{\prime}$, $\omega^{-}(\xi)=\emptyset$ by default. The set $\omega^{-}(\xi)$ is known as an $\alpha$-limit set of the original system $\dot{x}=f(x)$ (Hirsch et al. (2012)). We denote the set of all $\alpha$-limit sets in ${\mathcal{X}}$ by $\mathcal{W}^{-}$. For each $\Gamma\in\mathcal{W}^{-}$, its domain of attraction with respect to the time-reversed system is denoted by $D^{-}(\Gamma)$. Similar to $\mathcal{W}^{+}$, the set $\mathcal{W}^{-}$ contains all the equilibria and closed orbits in ${\mathcal{X}}$, and thus typically the intersection of $\mathcal{W}^{+}$ and $\mathcal{W}^{-}$ is not empty. Now we are ready to present the extension of Theorem 3.16:
Theorem 5.34
@@
    the set ${\mathcal{W}^{+}\cup\mathcal{W}^{-}}$ is finite or countable.
Then, the set $\{F(\Omega)\mid\Omega\in\mathcal{W}^{+}\cup\mathcal{W}^{-}\}$ has exactly one maximal element, that is, there exists $\Omega\in\mathcal{W}^{+}\cup\mathcal{W}^{-}$ such that $F(\Omega)\supseteq F(\Omega^{\prime})$ for all Ωsuperscriptnormal-$\Omega^{\prime}$ in $\mathcal{W}^{+}\cup\mathcal{W}^{-}$.
@@
    For every $\Omega^{+}\in\mathcal{W}^{+}$ and $\Omega^{-}\in\mathcal{W}^{-}$, there exist at least one precompact trajectory in $D^{+}(\Omega^{+})$ and one backward precompact trajectory in $D^{-}(\Gamma^{-})$,
@@
The proof of Theorem 5.34 is similar to that of Theorem 3.16, and can be found in Appendix B. Under conditions (T1’)-(T3’), Theorem 5.34 says that any continuous immersion $F$ cannot fully distinguish different limit sets in ${\mathcal{X}}$. Compared with Theorem 3.16, (T2’) is relaxed to allow diverging trajectories, as long as these trajectories converge to some limit sets (such as an unstable equilibrium) backward in time. At the same time, Theorem 5.34 requires that the time-reversed system $\dot{z}=-g(z)$ also has closed basins, which is trivially satisfied by any linear system.
Example 5.35
Consider the same system in Example 5.33 and ${\mathcal{X}}=\mathbb{R}^{2}\backslash\left((-1,0.5)\cup(0.5,1)\right)\times\{0\}$. Since (T2’)-(T4’) in Theorem 5.34 are satisfied, any linear immersion over ${\mathcal{X}}$ collapses the three equilibria into one. Therefore, no one-to-one linear immersions exist over this specific ${\mathcal{X}}$. Compared with the indirect extensions of Theorem 3.16, Theorem 5.34 is directly applied to show the non-existence of linear immersions, without constructing the subdomain ${\mathcal{X}}^{\prime}$ as in Example 5.33.
@@ 6 Conclusion
In this work, we first show that linear immersions collapse different $\omega$-limit sets into one under the condition that (i) all trajectories in ${\mathcal{X}}$ are precompact and (ii) there are at most countably many $\omega$-limit sets. Then we bridge our theoretical findings on exact linear immersions with approximate linear immersions learned from data. We show that as the size of the data set increases and the sampling interval decreases, the learned linear immersion converges to functions incapable of distinguishing different $\omega$-limit sets. To extend the applicability of our results beyond the constraints of precompact trajectories, we have also present several extensions to the main theorem. These extensions broaden the scope of our results, allowing us to address systems with diverging trajectories.
@@ Appendix A Proof of the Main Theorem
To prove Theorem 3.16, we first need to introduce two lemmas. The first lemma reveals a relation between $\omega$-limit sets of the original system and the immersed system.
Lemma A.36
@@
Remark A.37
Under condition (T2) in Theorem 3.16, for any $\Omega\in\mathcal{W}^{+}$, the image $\hat{\Omega}:=F(\Omega)$ is an $\omega$-limit set for the system $\dot{z}=g(z)$. Indeed, by definition there is some $\xi\in{\mathcal{X}}$ such that $\omega^{+}(\xi)=\Omega$. Thus, from Lemma A.36 we have that $\hat{\Omega}=F(\Omega)=F(\omega^{+}(\xi))=\omega^{+}(F(\xi))$.
@@
Proof. Since $\varphi(t,\xi)$ is precompact in ${\mathcal{X}}$, by Lemmas 2.10 and A.36, $\Omega:=\omega^{+}(\xi)$ is nonempty and $F(\Omega)=\omega^{+}(F(\xi))$. Let $x$ be a point in $D^{+}(\Omega)$. Then, there exists a sequence $t_{n}\geq 0$ such that $t_{n}\rightarrow+\infty$ and $\varphi(t_{n},x)\rightarrow x^{\prime}$ for some $x^{\prime}\in\Omega$. By (3) and the continuity of $F$, $\lim_{n\rightarrow+\infty}\psi(t_{n},F(x))=\lim_{n\rightarrow+\infty}F(\varphi(t_{n},x))=F(x^{\prime})\in F(\Omega)=\omega^{+}(F(\xi))$. Hence, $F(x)\in D^{+}(\omega^{+}(F(\xi)))$ and thereby $F(D^{+}(\omega^{+}(\xi)))\subseteq D^{+}(F(\omega^{+}(\xi)))$. $\Box$
Proof of Theorem 3.16: Since by (T2) every trajectory is precompact in ${\mathcal{X}}$ and by (T3) there are at most countably many $\omega$-limit sets in ${\mathcal{X}}$, we have that ${\mathcal{X}}=\bigcup_{i\in I}D^{+}(\Omega_{i})$, for a finite or countable set $I$. Thus, $F({\mathcal{X}})=\bigcup_{i\in I}F(D^{+}(\Omega_{i}))$. By (T2) and Lemmas A.36 and A.38, $F(\Omega_{i})$ is an $\omega$-limit set for all $i\in I$ and
@@
According (T1), $D^{+}(F(\Omega_{i}))$ are closed in ${\mathcal{Z}}$ and thus $D^{+}(F(\Omega_{i}))\cap F({\mathcal{X}})$ are closed in the subspace topology induced on $F({\mathcal{X}})$ and disjoint for all $i\in I$. That is, $F({\mathcal{X}})$ is a disjoint union of a countable collection of closed sets. Since ${\mathcal{X}}$ is path-connected and $F$ is continuous, $F({\mathcal{X}})$ is path-connected as well. Thus, by a theorem of Sierpiński (1918), only one of the sets in the collection $\{D^{+}(F(\Omega_{i}))\cap F({\mathcal{X}})\}_{i\in I}$ can be nonempty. However, since points in $F(\Omega_{i})$ are limit points of $D^{+}(F(\Omega_{i}))$ and $D^{+}(F(\Omega_{i}))$ is closed, the set $D^{+}(F(\Omega_{i}))\cap F({\mathcal{X}})$ contains the nonempty set $F(\Omega_{i})$ for all $i\in I$. Therefore, $D^{+}(F(\Omega_{i}))\cap F({\mathcal{X}})$ must be equal to $F({\mathcal{X}})$ for all $i\in I$.
Furthermore, $F(\Omega_{i})$ must be the same for all $i\in I$ since their domains of attraction intersect each other. $\Box$
Remark A.39
Sierpiński’s Theorem states that if a continuum ${\mathcal{X}}$ has a countable cover $\{X_{i}\}_{i=1}^{\infty}$ by pairwise disjoint closed subsets, then at most one of the sets $X_{i}$ is non-empty. A continuum is a compact connected Hausdorff space, but we do not assume that ${\mathcal{X}}$ is compact. However, the theorem is still true if ${\mathcal{X}}$ is not compact. Indeed, suppose that two of the sets $X_{i}$ would be nonempty, and pick two points $p,q$, one in each set. Consider a (continuous) path $\gamma:[0,1]\rightarrow{\mathcal{X}}$ that joins these two points, and let $\Gamma:=\gamma([0,1])$. Now the sets $\{X_{i}\bigcap\Gamma\}_{i=1}^{\infty}$ form a disjoint cover of the continuum $\Gamma$, but two of these sets are nonempty, a contradiction.
@@
We first show a property of systems with closed basins.
Lemma B.40
@@
Thus, we have $\Omega\subseteq D^{-}(\Gamma)$ and $\Gamma\subseteq D^{+}(\Omega)$. Next, it can be shown that any limit set is closed and invariant in ${\mathcal{X}}$ (Alligood et al. (1996)). Since $\Omega$ is invariant and $\Omega\subseteq D^{-}(\Gamma)$, we have $\Gamma\subseteq\text{cl}(\Omega)=\Omega$. Similarly, we have $\Omega\subseteq\text{cl}(\Gamma)=\Gamma$. Thus, $\Omega=\Gamma$. $\Box$
Next, we extend Lemmas A.36 and A.38 for $\alpha$-limit sets.
Lemma B.41
Let $F$ be an immersion that maps $\dot{x}=f(x)$ on ${\mathcal{X}}$ to $\dot{z}=g(x)$ on ${\mathcal{Z}}$. For any $\xi\in{\mathcal{X}}$, if $\omega^{-}(\xi)$ exists, then $\omega^{-}(F(\xi))$ exists and contains $F(\omega^{-}(\xi))$. Furthermore, if the trajectory starting at $\xi$ is backward precompact in ${\mathcal{X}}$, ${F(\omega^{-}(\xi))=\omega^{-}(F(\xi))}$.
@@
Lemma B.42
Let $F$ be an immersion that maps $\dot{x}=f(x)$ on ${\mathcal{X}}$ to a system $\dot{z}=g(z)$ on ${\mathcal{Z}}$. For any $\xi\in{\mathcal{X}}$, if $\varphi(t,\xi)$ is precompact in $X$, then
@@
The inclusion relation when $\varphi(t,\xi)$ is precompact in ${\mathcal{X}}$ is directly implied by Lemma A.38 . The inclusion relation when $\varphi(t,\xi)$ is backward precompact in ${\mathcal{X}}$ can be shown via similar arguments in the proof of Lemma A.38 with Lemma 2.10, Lemma B.41, and (44). $\Box$
Proof of Theorem 5.34: Let $\widehat{\mathcal{W}^{+}}$ be the set of $\omega$-limit sets $\Omega$ in $\mathcal{W}^{+}$ whose domain of attraction $D^{+}(\Omega)$ contains at least one trajectory precompact in ${\mathcal{X}}$. Similarly, we define $\widehat{\mathcal{W}^{-}}$ be the set of $\alpha$-limit sets $\Gamma$ in $\mathcal{W}^{-}$ whose domain of attraction $D^{-}(\Gamma)$ contains at least one trajectory backward precompact in ${\mathcal{X}}$. Then, for any limit set $\Omega\in\widehat{\mathcal{W}^{+}}\cup\widehat{\mathcal{W}^{-}}$, we define

### v5

@@ 5 Extensions of the Main Theorem
Due to the condition (T2), Theorem 1 cannot be directly applied for systems with diverging trajectories in $\mathcal{X}$. In this section, we show several extensions to Theorem 1 that accommodate systems with diverging trajectories.
5.1 Indirect Extensions of Theorem 1
The following two propositions, in conjunction with Theorem 1, show the non-existence of one-to-one linear immersions even when a diverging trajectory is present. We refer to these propositions as “indirect extensions” to Theorem 1 because they must be utilized in conjunction with it.
Proposition 5.6.
@@ 5.1 Indirect Extensions of Theorem 1
Proposition 5.7.
Suppose that $\mathcal{X}^{\prime}\subseteq\mathcal{X}$ is forward invariant for the time-reversed system $\dot{x}=-f(x)$. If $F$ immerses the system $\dot{x}=f(x)$ over $\mathcal{X}$ into $\dot{z}=g(z)$ over $\mathcal{Z}$, then the same $F$ also immerses the time-reversed system $\dot{x}=-f(x)$ over $\mathcal{X}^{\prime}$ into $\dot{z}=-g(z)$ over $F(\mathcal{X}^{\prime})$.
@@
Thus, the solution $\psi(-t,F(\xi))$ of $\dot{z}=-g(z)$ is well defined and equal to $F(\varphi(-t,\xi))$ for all $t\geq 0$. $\Box$ By Proposition 5.7, if we can show that a one-to-one linear immersion does not exist for the time-reversed system $\dot{x}=-f(x)$ over a forward-invariant subdomain of $\mathcal{X}$, then there is no one-to-one linear immersion for the original system $\dot{x}=f(x)$ over $\mathcal{X}$. The following example demonstrates how to extend our results to systems with diverging trajectories by combining the above two propositions with Theorem 1.
@@
Consider a system in $\mathbb{R}^{2}$ with the phase portrait shown in Fig. 6(a). The system has three $\omega$-limit sets $\{-1,0,1\}$, where $-1$ and $1$ are unstable equilibria and 00 is a saddle point. We want to show that there is no one-to-one linear immersion $F$ of this system over $\mathcal{X}=\mathbb{R}^{2}$. To achieve this goal, we cannot directly apply Corollary 3 since any trajectory starting outside $[-1,1]\times\{0\}$ is not precompact. However, by applying Theorem 1 to the forward-invariant subdomain $\mathcal{X}^{\prime}=[-1,0]\times\{0\}$ (or $[0,1]\times\{0\}$) of $\mathcal{X}$, we know that there is no one-to-one linear immersion over $\mathcal{X}^{\prime}$. Then, by Proposition 5.6, no one-to-one linear immersion exists over $\mathcal{X}$.
@@
To make this example more challenging, consider $\mathcal{X}=\mathbb{R}^{2}\backslash\left((-1,-0.5)\cup(0.5,1)\right)\times\{0\}$. Note that our previous argument does not work anymore since any forward invariant $\mathcal{X}^{\prime}\subseteq\mathcal{X}$ contains at least one trajectory that is not precompact. Now consider the time-reversed system, shown by the phase portrait in Fig. 6(b). Since $\mathcal{X}$ is not forward invariant for the time-reversed system, we take the forward invariant subset $\mathcal{X}^{\prime}=\mathbb{R}^{2}\backslash\left((-1,0)\cup(0,1)\right)\times\{0\}$ of $\mathcal{X}$. Since all trajectories of the time-reversed systems in $\mathcal{X}^{\prime}$ are precompact, Corollary 3 implies that there is no one-to-one linear immersion for the time-reversed system over $\mathcal{X}^{\prime}$. Then, by Proposition 5.7, we know there is no one-to-one linear immersion for the original system over $\mathcal{X}$.
5.2 A Direct Extension of Theorem 1
@@
While Propositions 5.6 and 5.7 allow us to apply Theorem 1 to systems with diverging trajectories, identifying the appropriate subdomain $\mathcal{X}^{\prime}$ for more complex examples can be challenging. In this section, we present a direct extension to Theorem 1, where we replace the original condition (T2) with a weaker condition. Before delving into this extension, we introduce several key definitions.
@@
For a system $\dot{x}=f(x)$ defined over $\mathcal{X}$, let $\mathcal{X}^{\prime}$ be the set of states $\xi\in\mathcal{X}$ such that $\varphi(-t,\xi)$ is defined and contained in $\mathcal{X}$ for all $t\geq 0$ (Namely $\mathcal{X}^{\prime}$ is the largest backward invariant subset of $\mathcal{X}$ with respect to the system). Given an initial state $\xi\in\mathcal{X}$, the trajectory $\varphi(t,\xi)$ from $\xi$ is called backward precompact in $\mathcal{X}$ if $\xi\in\mathcal{X}^{\prime}$ and the trajectory from $\xi$ is precompact in $\mathcal{X}$ with respect to the time-reversed system $\dot{x}=-f(x)$.
@@
For each initial state $\xi\in\mathcal{X}$, the $\omega$-limit set of $\xi$ in $\mathcal{X}$ with respect to the time-reversed system $\dot{x}=-f(x)$ is denoted by $\omega^{-}(\xi)$. If $\xi\not\in\mathcal{X}^{\prime}$, we define $\omega^{-}(\xi)=\emptyset$. The set $\omega^{-}(\xi)$ is known as an $\alpha$-limit set of the original system $\dot{x}=f(x)$ [12]. For any subset $\Gamma$ of $\mathcal{X}$, its domain of attraction with respect to the time-reversed system is denoted by $D^{-}(\Gamma)$. We denote the set of all nonempty $\alpha$-limit sets in $\mathcal{X}$ by $\mathcal{W}^{-}$. Similar to $\mathcal{W}^{+}$, the set $\mathcal{W}^{-}$ contains all the equilibria and closed orbits in $\mathcal{X}$, and thus typically the intersection of $\mathcal{W}^{+}$ and $\mathcal{W}^{-}$ is not empty. Now we are ready to present the extension of Theorem 1:
Theorem 5.9.
@@
    the set ${\mathcal{W}^{+}\cup\mathcal{W}^{-}}$ is countable.
Then, the set $\{F(\Omega)\mid\Omega\in\mathcal{W}^{+}\cup\mathcal{W}^{-}\}$ has exactly one maximal element, that is, there exists $\Omega\in\mathcal{W}^{+}\cup\mathcal{W}^{-}$ such that $F(\Omega)\supseteq F(\Omega^{\prime})$ for all $\Omega^{\prime}$ in $\mathcal{W}^{+}\cup\mathcal{W}^{-}$.
@@
    For every $\Omega^{+}\in\mathcal{W}^{+}$ and $\Omega^{-}\in\mathcal{W}^{-}$, there exist at least one precompact trajectory in $D^{+}(\Omega^{+})$ and one backward precompact trajectory in $D^{-}(\Omega^{-})$,
@@
The proof of Theorem 5.9 is similar to that of Theorem 1, and can be found in Appendix B. Under conditions (T1’)-(T3’), Theorem 5.9 says that any continuous immersion $F$ cannot fully distinguish different limit sets in $\mathcal{X}$. Compared with Theorem 1, (T2’) is relaxed to allow diverging trajectories, as long as these trajectories converge to some limit sets (such as an unstable equilibrium) backward in time. At the same time, Theorem 5.9 requires that the time-reversed system $\dot{z}=-g(z)$ also has closed basins, which is trivially satisfied by any linear system.
Example 5.10.
Consider the same system in Example 5.8 and $\mathcal{X}=\mathbb{R}^{2}\backslash\left((-1,-0.5)\cup(0.5,1)\right)\times\{0\}$. Since (T2’)-(T4’) in Theorem 5.9 are satisfied, any linear immersion over $\mathcal{X}$ collapses the three equilibria into one. Therefore, no one-to-one linear immersions exist over this specific $\mathcal{X}$. Compared with the indirect extensions of Theorem 1, Theorem 5.9 is directly applied to show the non-existence of linear immersions, without constructing the subdomain $\mathcal{X}^{\prime}$ as in Example 5.8.
@@ 6 Conclusion
In this work, we first show that linear immersions collapse different $\omega$-limit sets into one under the condition that (i) all trajectories in $\mathcal{X}$ are precompact and (ii) there are at most countably many $\omega$-limit sets. Then we bridge our theoretical findings on exact linear immersions with approximate linear immersions learned from data. We show that as the size of the data set increases and the sampling interval decreases, the learned linear immersion converges to functions incapable of distinguishing different $\omega$-limit sets. To extend the applicability of our results beyond the constraints of precompact trajectories, we have also presented several extensions to the main theorem. These extensions broaden the scope of our results, allowing us to address systems with diverging trajectories. For future work, we are interested in the effect of multiple $\omega$-limit sets on the approximation error of Koopman-based projected models [11], which is well-defined even if a linear immersion does not exist. {ack} We thank Matthew Kvalheim for helpful discussions about the relation of our results with those in [16].
@@ Appendix A Proof of the Main Theorem
To prove Theorem 1, we first need to introduce two lemmas. Recall that by Remark 2, an immersion is assumed to be continuous unless otherwise specified. The first lemma reveals a relation between $\omega$-limit sets of the original system and the immersed system.
Lemma A.11.
@@
Remark A.12.
Under condition (T2) in Theorem 1, for any $\Omega\in\mathcal{W}^{+}$, the image $\hat{\Omega}:=F(\Omega)$ is an $\omega$-limit set for the system $\dot{z}=g(z)$. Indeed, by definition there is some $\xi\in\mathcal{X}$ such that $\omega^{+}(\xi)=\Omega$. Thus, from Lemma A.11 we have that $\hat{\Omega}=F(\Omega)=F(\omega^{+}(\xi))=\omega^{+}(F(\xi))$.
@@
Proof. Since $\varphi(t,\xi)$ is precompact in $\mathcal{X}$, by Lemmas 1 and A.11, $\Omega:=\omega^{+}(\xi)$ is nonempty and $F(\Omega)=\omega^{+}(F(\xi))$. Let $x$ be a point in $D^{+}(\Omega)$. Then, there exists a sequence $t_{n}\geq 0$ such that $t_{n}\rightarrow+\infty$ and $\varphi(t_{n},x)\rightarrow x^{\prime}$ for some $x^{\prime}\in\Omega$. By (3) and the continuity of $F$, $\lim_{n\rightarrow+\infty}\psi(t_{n},F(x))=\lim_{n\rightarrow+\infty}F(\varphi(t_{n},x))=F(x^{\prime})\in F(\Omega)=\omega^{+}(F(\xi))$. Hence, $F(x)\in D^{+}(\omega^{+}(F(\xi)))$ and thereby $F(D^{+}(\omega^{+}(\xi)))\subseteq D^{+}(F(\omega^{+}(\xi)))$. $\Box$
Proof of Theorem 1: Since by (T2) every trajectory is precompact in $\mathcal{X}$ and by (T3) there are at most countably many $\omega$-limit sets in $\mathcal{X}$, we have that $\mathcal{X}=\bigcup_{i\in I}D^{+}(\Omega_{i})$, for a countable set $I$. Thus, $F(\mathcal{X})=\bigcup_{i\in I}F(D^{+}(\Omega_{i}))$. By (T2) and Lemmas A.11 and A.13, $F(\Omega_{i})$ is an $\omega$-limit set for all $i\in I$ and
@@
According (T1), $D^{+}(F(\Omega_{i}))$ is closed in $\mathcal{Z}$ and thus $D^{+}(F(\Omega_{i}))\cap F(\mathcal{X})$ is closed in the subspace topology induced on $F(\mathcal{X})$ for all $i\in I$. Moreover, $D^{+}(F(\Omega_{i}))\cap F(\mathcal{X})$ for all $i\in I$ is nonempty since points in $F(\Omega_{i})$ are limit points of $D^{+}(F(\Omega_{i}))$ and thus contained in the closed set $D^{+}(F(\Omega_{i}))\cap F(\mathcal{X})$.
Finally, since the domains of attraction of two different $\omega$-limit sets must be disjoint, for any indices $i$ and $i^{\prime}\in I$, we have either (a) the two $\omega$-limit sets $F(\Omega_{i})$ and $F(\Omega_{i^{\prime}})$ of the immersed system are equal, which implies $D^{+}(F(\Omega_{i}))=D^{+}(F(\Omega_{i^{\prime}}))$; or (b) $D^{+}(F(\Omega_{i}))$ and $D^{+}(F(\Omega_{i^{\prime}}))$ are disjoint.
Suppose there exist $i$ and $i^{\prime}$ such that case (b) holds, that is, $D^{+}(F(\Omega_{i}))$ and $D^{+}(F(\Omega_{i^{\prime}}))$ are disjoint. It follows from (36) that $F(\mathcal{X})$ is a disjoint union of a countable collection of closed sets. Since $\mathcal{X}$ is path-connected and $F$ is continuous, $F(\mathcal{X})$ is path-connected as well. Thus, by the main theorem of [26], only one of the sets in the collection $\{D^{+}(F(\Omega_{i}))\cap F(\mathcal{X})\}_{i\in I}$ can be nonempty. This contradicts our earlier proof that $D^{+}(F(\Omega_{i}))\cap F(\mathcal{X})$ is nonempty for all $i\in I$. Thus by contradiction, $F(\Omega_{i})$ must be the same for all $i\in I$. $\Box$
Remark A.14.
Sierpiński’s Theorem states that if a continuum $\mathcal{X}$ has a countable cover $\{X_{i}\}_{i=1}^{\infty}$ by pairwise disjoint closed subsets, then at most one of the sets $X_{i}$ is non-empty. A continuum is a compact connected Hausdorff space, but we do not assume that $\mathcal{X}$ is compact. However, the theorem is still true if $\mathcal{X}$ is not compact. Indeed, suppose that two of the sets $X_{i}$ would be nonempty, and pick two points $p,q$, one in each set. Consider a (continuous) path $\gamma:[0,1]\rightarrow\mathcal{X}$ that joins these two points, and let $\Gamma:=\gamma([0,1])$. Now the sets $\{X_{i}\cap\Gamma\}_{i=1}^{\infty}$ form a disjoint cover of the continuum $\Gamma$, but two of these sets are nonempty, a contradiction.
@@
Recall that the definitions of backward precompact trajectories, $\alpha$-limit sets, and their domains of attraction are provided in Section 5.2. We first show a property of systems with closed basins.
Lemma B.15.
@@
Thus, we have $\Omega\subseteq D^{-}(\Gamma)$ and $\Gamma\subseteq D^{+}(\Omega)$. Next, it can be shown that any limit set is closed and invariant in $\mathcal{X}$ [1]. Since $\Omega$ is invariant and $\Omega\subseteq D^{-}(\Gamma)$, we have $\Gamma\subseteq\text{cl}(\Omega)=\Omega$. Similarly, we have $\Omega\subseteq\text{cl}(\Gamma)=\Gamma$. Thus, $\Omega=\Gamma$. $\Box$
Next, we extend Lemmas A.11 and A.13 for $\alpha$-limit sets.
Lemma B.16.
Let $F$ be an immersion that maps $\dot{x}=f(x)$ on $\mathcal{X}$ to $\dot{z}=g(z)$ on $\mathcal{Z}$. For any $\xi\in\mathcal{X}$, if $\omega^{-}(\xi)$ exists, then $\omega^{-}(F(\xi))$ exists and contains $F(\omega^{-}(\xi))$. Furthermore, if the trajectory starting at $\xi$ is backward precompact in $\mathcal{X}$, ${F(\omega^{-}(\xi))=\omega^{-}(F(\xi))}$.
@@
Lemma B.17.
Let $F$ be an immersion that maps $\dot{x}=f(x)$ on $\mathcal{X}$ to a system $\dot{z}=g(z)$ on $\mathcal{Z}$. For any $\xi\in\mathcal{X}$, if $\varphi(t,\xi)$ is precompact in $\mathcal{X}$, then
@@
The inclusion relation when $\varphi(t,\xi)$ is precompact in $\mathcal{X}$ is directly implied by Lemma A.13 . The inclusion relation when $\varphi(t,\xi)$ is backward precompact in $\mathcal{X}$ can be shown via similar arguments in the proof of Lemma A.13 with Lemma 1, Lemma B.16, and (35). $\Box$
The following lemma is a key component in the proof of Theorem 5.9.
Lemma B.18.
Let $\widehat{\mathcal{W}}^{+}$ be the set of $\omega$-limit sets $\Omega$ in $\mathcal{W}^{+}$ whose domain of attraction $D^{+}(\Omega)$ contains at least one trajectory precompact in $\mathcal{X}$, and let $\widehat{\mathcal{W}}^{-}$ be the set of $\alpha$-limit sets $\Gamma$ in $\mathcal{W}^{-}$ whose domain of attraction $D^{-}(\Gamma)$ contains at least one trajectory backward precompact in $\mathcal{X}$. Then, under (T1’)-(T3’), the images $F(\Omega)$ and $F(\Omega^{\prime})$ are the same for any $\Omega$ and $\Omega^{\prime}$ in $\widehat{\mathcal{W}}^{+}\cup\widehat{\mathcal{W}}^{-}$.
{pf}
For any limit set $\Omega\in\widehat{\mathcal{W}}^{+}\cup\widehat{\mathcal{W}}^{-}$, let us define